Which is the equation of a parabola with focus (0, 5) and directrix y= -5?
Answer by lwsshak3(11628)

Which is the equation of a parabola with focus (0, 5) and directrix y= -5?
***
Given parabola opens upwards.
Its basic form of equation:
(x-h)^2=4p(y-k), (h,k)=coordinates of the vertex
vertex: (0,0)
axis of symmetry: x=0
p=5(distance from vertex to focus or directrix on the axis of symmetry)
4p=20
equation: x^2=20y